If a-b=15 and a^2+b^2=137, find the numerical ab.
Answer by Jk22(389)

(a-b)^2=a^2-2ab+b^2=15*15=150+50+25=225

hence a^2+b^2-(a-b)^2=2ab=137-225=-(63+25)=-88

=> ab=-44